# Approximate $SU(5)$, Fine Structure Constants

### v1

@@ 11.4 Table of Combinations
The most important outcome of the fluctuating-size-of-links lattice, we propose, is that it gives us te possibility of having a Planck scale very different from te “unification scale” and still claim a “fundamental” lattice at the unification scale. But we would of course like to see, if the order of magnitudes are at all thinkable. We therefore in figure LABEL:ill illustrate how we imagine a smooth Gaussian distribution in the logarith of the link length say.
@@ 11.7 Fitted $\mu_{0}$
The choice of $\mu_{0}$ that would make prediction perfect was as seen from the table and our calculation $24TeV$, which is higher than the top quark mass only by a factor $\frac{24.23}{172.25GeV}=7.11$. Very speculatively one could attempt to construct some fitting of the density as a function of the scale fermions still findable as effectively massles at the scales considered. Above the top mass of course the effective number of massless Fermions at the scale correspond to 3 families, but as one asks below the top qaurk mass there is a 1/3 or 1/4 of a family missing, and one could claim that just below the top mass we have some 3-1/4 or 3-1/3 families left, and then crudely estimate in this spirit, if one could fit the (non-integer) number of families to a function reaching a bit above the top-mass a maximum of 3 families. Then this maximum in a curve taken as a function of the logartihm of the scale would have its maximum with value 3 families very close to our wanted $\mu_{0}=24.23TeV.$ (most wellcome to make gravity scale match our model).
@@
On a logarithmic mass scale there is a “density of species” of 1 over a scale dististance $\ln\frac{m_{t}}{m_{b}}=6.02$, meaning density = 0.166 per e-factor. The center of this interval is the geometric mean of 172.25 GeV and 4.180 GeV, which is 26.83 GeV. Next take the interval beteen the b-mass and the s-mass which has length $\ln(\frac{4.180}{0.095})=3.784$ and contains two species quarks and 3 species fermions, if we include the $\tau$-lepton. This means for this region around $\sqrt{4.180*0.095}GeV$ = 0.630 GeV we have the density of quark species $2/3.784=0.5285$. This density is 3.184 times bigger than in the interval between t and b. If the deviation from the maximal number 3 of the number of families at the scale being seen as massless were varying with the logarithm of the scale quadratically counted out from some center value of the scale, then the slope of it, being in fact the “density of species” would vary linearly, and we should just extrapolate the density to the point, where it passes zero to find the maximum point for the formal number of massless families. In logarithmthe distance between the two points we considered is $\ln(\frac{26.83}{0.630})=3.75$,and the linear extrapolation leads to the zero for the slope point displaced upward from the 26.83 by the exponential of 3.75/(3.184-1)= 1.717, and that gives 149.4 GeV. It is actually very close to the top-mass. Now we see, that if we take the difference between the two possibilities we mention in the table for the Planck scale as an estimate of the uncertainty of only order of magnitude numbers, then this uncertainty for the ratio given in the table is of order 0.06. But the best of the points actually for the top mass as $\mu_{0}$ deviates only by 0.03 from the predicted 1.5, so we must say that we shouls take it as agreement within expected uncertainty.
@@ 12.3 Gravity has to be “Weak” on Fundamental Scale
This may bring us some message about gravity: We have to invent a story of the kind, that gravity is for some reason very weak compared to the fundamental scale expectation. Our above described model namely has as its philosophy, that the unified scale - which remains low compared to the Planck scale in energy - is to be the “ fundamental scale”! You might speculatively think about, that the $g^{\mu\nu}$ (with upper indices) has appeared as kind of spontaneous breaking of e.g. diffeomorphism symmetry, and and thus has a chance to be small (often one finds reltively small spontaneously breaking fields, otherwise it would not be so common with low temperature super conductivity, that it was a big sensation to find high temperature super conductivity). If this $g^{\mu\nu}$ is small compared the our fundamental lattice, then compared to this lattice the $g_{\mu\nu}$ with lower indices will be large and thus the length say of a lattice link would be big. This bigness would be bigness compared to the Planck constant and so getting $g^{\mu\nu}$ by some spontaneous breaking story would help bringing about the lack of coincidence of our fundamental scale with the Planck one[22].
Although this idea of having $g^{\mu\nu}$ representing a smontaneous symmetry break down and being “small” for that reason, seems attrcktive to me, we shall in this article rather seek to solve the problem with the Planck scale being different from “the our unified one” $\mu_{U}$ by the idea of fluctuating lattice link size described in next subsection.
@@ Acknowledgement
I remember, that Svend Erik Rugh was the first to tell me, that in fact the SU(5) GUT coupling was crittical (for SUSY-GUT) the unified coupling inverted is down to $1/\alpha_{5}(10^{16}GeV)\sim 25$ and closer to just one of the above mentioned critical (242), $15.3\pm 7.5\%$, but without the factor 3.

### v3

@@ 11.4 Table of Combinations
The most important outcome of the fluctuating-size-of-links lattice, we propose, is that it gives us the possibility of having a Planck scale very different from the “unification scale” and still claim a “fundamental” lattice at the unification scale. But we would of course like to see, if the order of magnitudes are at all thinkable. We therefore in figure 11.4illustrate how we imagine a smooth Gaussian distribution in the logarith of the link length say.
@@ 11.7 Fitted $\mu_{0}$
The choice of $\mu_{0}$ that would make prediction perfect was as seen from the table and our calculation $24TeV$, which is higher than the top quark mass only by a factor $\frac{24.23}{172.25GeV}=7.11$. Very speculatively one could attempt to construct some fitting of the density as a function of the scale fermions still findable as effectively massless at the scales considered. Above the top mass of course the effective number of massless Fermions at the scale correspond to 3 families, but as one asks below the top qaurk mass there is a 1/3 or 1/4 of a family missing, and one could claim that just below the top mass we have some 3-1/4 or 3-1/3 families left, and then crudely estimate in this spirit, if one could fit the (non-integer) number of families to a function reaching a bit above the top-mass a maximum of 3 families. Then this maximum in a curve taken as a function of the logartihm of the scale would have its maximum with value 3 families very close to our wanted $\mu_{0}=24.23TeV.$ (most wellcome to make gravity scale match our model).
@@
On a logarithmic mass scale there is a “density of species” of 1 over a scale dististance $\ln\frac{m_{t}}{m_{b}}=6.02$, meaning density = 0.166 per e-factor. The center of this interval is the geometric mean of 172.25 GeV and 4.180 GeV, which is 26.83 GeV. Next take the interval beteen the b-mass and the s-mass which has length $\ln(\frac{4.180}{0.095})=3.784$ and contains two species quarks and 3 species fermions, if we include the $\tau$-lepton. This means for this region around $\sqrt{4.180*0.095}GeV$ = 0.630 GeV we have the density of quark species $2/3.784=0.5285$. This density is 3.184 times bigger than in the interval between t and b. If the deviation from the maximal number 3 of the number of families at the scale being seen as massless were varying with the logarithm of the scale quadratically counted out from some center value of the scale, then the slope of it, being in fact the “density of species” would vary linearly, and we should just extrapolate the density to the point, where it passes zero to find the maximum point for the formal number of massless families. In logarithmthe distance between the two points we considered is $\ln(\frac{26.83}{0.630})=3.75$,and the linear extrapolation leads to the zero for the slope point displaced upward from the 26.83 by the exponential of 3.75/(3.184-1)= 1.717, and that gives 149.4 GeV. It is actually very close to the top-mass. Now we see, that if we take the difference between the two possibilities we mention in the table for the Planck scale as an estimate of the uncertainty of only order of magnitude numbers, then this uncertainty for the ratio given in the table is of order 0.06. But the best of the points actually for the top mass as $\mu_{0}$ deviates only by 0.03 from the predicted 1.5, so we must say that we should take it as agreement within expected uncertainty.
@@ 12.2 Problem with Planck Scale in Our Model
(In the references [16, 15] we have after the present made progress on the fluctuating lattice, meant help on the problemof unified and Planck scale not matching)
@@ 12.3 Gravity Has to Be “Weak” on Fundamental Scale
This may bring us some message about gravity: We have to invent a story of the kind, that gravity is for some reason very weak compared to the fundamental scale expectation. Our above described model namely has as its philosophy, that the unified scale - which remains low compared to the Planck scale in energy - is to be the “ fundamental scale”! You might speculatively think about, that the $g^{\mu\nu}$ (with upper indices) has appeared as kind of spontaneous breaking of e.g. diffeomorphism symmetry, and and thus has a chance to be small (often one finds relatively small spontaneously breaking fields, otherwise it would not be so common with low temperature super conductivity, that it was a big sensation to find high temperature super conductivity). If this $g^{\mu\nu}$ is small compared the our fundamental lattice, then compared to this lattice the $g_{\mu\nu}$ with lower indices will be large and thus the length say of a lattice link would be big. This bigness would be bigness compared to the Planck constant and so getting $g^{\mu\nu}$ by some spontaneous breaking story would help bringing about the lack of coincidence of our fundamental scale with the Planck one[24].
Although this idea of having $g^{\mu\nu}$ representing a spontaneous symmetry break down and being “small” for that reason, seems attracktive to me, we shall in this article rather seek to solve the problem with the Planck scale being different from “the our unified one” $\mu_{U}$ by the idea of fluctuating lattice link size described in next subsection.
@@ 12.5 Baryon Non-Conservation ?
When we in this way have no other breaking of the baryon number than via the instantons present in the Standard Model, it means that our model points towards baryon assymetry should be caused via a lepton assymetry. We namely only get the instanton baryon number variation to wash away any previous baryon number assymetry unless there is a lepton number assymmetry. But this pointing to the baryon assymetry comming from a lepton assymetry is not an absolute must, because the baryon number concervation is still only an accidental symmetry in our model; we only got rid of the for true SU(5) so embarrasing violation of baryon conservation.
That we have a see-saw scale much lower than the Planck scale may also encourage us that even baryon violation could be there too, if it could help the assymetry without making decays. But we do favor leptonassymetry as the mechanism.
@@ 12.8.3 Small Hierarchy by the Charges from $G_{SMG}\times\cdots\times G_{SMG}$
12.8.4 Modification of Accurate Standard Model Results
The idea of the fluctuating lattice put forward above and the hypotesis that the lattice is truly existing means that ideal Standard Model perturbative calcuæations studied with high accuracy, such as e.g. the famous anomalous magnetic moment calculations, would be modified because they should be performed with truly exisitng lattice, which only has a finite link-length. In fact in the philosophy of the fluctuating lattice stricly speaking the central value for the link distribution is about the top mass or better say $10^{4}GeV$, which means that corrections to ideal perturbation theory would be much closer than if the most fundamental scale had been the Planck scale.
Also it would be very great of course, if one could experimentally somehow explore the non-locality due to relatively seldom lattice links of exceptionally large lengths. Indeed also such effects should be much more accesible if the present fluctuating lattice was true, than say Planck scale were the fundamental scale.
@@ Acknowledgement
I remember, that Svend Erik Rugh was the first to tell me, that in fact the SU(5) GUT coupling was crittical (for SUSY-GUT) the unified coupling inverted is down to $1/\alpha_{5}(10^{16}GeV)\sim 25$ and closer to just one of the above mentioned critical (242), $15.3\pm 7.5\%$, but without the factor 3. I thenk the referees for small additions, e.g subsections 1.1 and 12.8.4.